The perimeter of a train ticket is 46 centimeters. The area is 112 square centimeters. What are the dimensions of the ticket?
Answer by linht117(37)

112/16= 7
length= 16
width= 7
check
p= 2(l+w)
46=2(16+7)
46=46
a=l*w
112=16*7
112=112